How can I find the sum of the first n multiples of 4?
Answer by fractalier(6550)

The sum of an arithmetic series is found by
Sn = (n/2)(a1 + an)
Here an = 4n and a1 = 4, so
Sn = (n/2)(4 + 4n) = n(2 + 2n) = 2n^2 + 2n